OpenStudy (princesssleelee):

What is the value of 373 using the identity (x − y)3 = x3 − 3x2y + 3xy2 − y3? Hint: 373 = (40 − 3)3; therefore, x = 40 and y = 3.

OpenStudy (princesssleelee):

Overally plugging in the x and y equaling : 37^3=(40-3)^3=40^3-3*40^2*3+3*40*3^2-3^3=

OpenStudy (princesssleelee):

I do not know what it would equal overall :(

OpenStudy (princesssleelee):

@Mehek14

Directrix (directrix):

Identity (x − y)^3 = x^3 − 3x^2y + 3xy^2 − y^3 373 = (40 − 3)^3 Replace x by 40 and y by 3 37^3 = (40 − 3)^3 = x^3 − 3x^2y + 3xy^2 − y^3

Directrix (directrix):

37^3 = (40 − 3)^3 = 40^3 − 3*40^2 *3 + 3*40 * 3^2 − 3^3 Crank out this: 40^3 − 3*40^2 *3 + 3*40 * 3^2 − 3^3

Directrix (directrix):

40^3 = − 3*40^2 *3 = + 3*40 * 3^2 = − 3^3 = Get the numerical value of each term and then add those.

OpenStudy (princesssleelee):

i got -6947 @Directrix

OpenStudy (princesssleelee):

Im not sure if that seems correct, however I just may be over thinking

OpenStudy (princesssleelee):

@pooja195

Directrix (directrix):

Let me see what I get.

Directrix (directrix):

40^3 = 64 000 − 3*40^2 *3 = -14 400 3*40 * 3^2 = 1 080 − 3^3 = - 27 ---------------------------------- We have to add all this.

Directrix (directrix):

>>i got -6947 I did not get that.

OpenStudy (princesssleelee):

I got that as well lol

Directrix (directrix):

I got this after I added the four parts: 50 653

Directrix (directrix):

If this is a "show your work" problem, you'll need to include all of the calculations from the expansion in your work.

OpenStudy (princesssleelee):

oh.... i added them and got the -6947, let me try again

OpenStudy (princesssleelee):

I am still adding them and getting my answer... @Directrix

Directrix (directrix):

Add two at the time. Then, add that sum to the next one.

Directrix (directrix):

I think you may be missing this one: 40^3 = 64 000